sue invested 19800in two mutual accounts one earned 6% and one earned 10% at the end of the year the total amount of interest earned was 1.492 how much did she invest in each fund
Answer by mananth(12270)

let her invest $ x in mutual fund A
so she will invest 19800-x in mutual fund B
.
Interest earned on eache = total interest
0.06x+0.1(19800-x)=1492
0.06x + 1980-0.1x=1492
-0.04x =1492-1980
x=-488/-0.04
=12200
She invested $12200 in Mutual fund A and the balance in mutual fund B